OpenStudy (anonymous):

HELP!! Use substitution to solve the system x − 5y = −27 (1) 7x − 6y = −15 (2) a) Find the value of y. (Answer with −100 if there is no solution.)

OpenStudy (anonymous):

x=5y-27 7(5y-27)-6y=-15 35y-189-6y=-15 29y=174 y=6 x=5(6)-27 x=3 (3,6) is the point that satisfies your system.

OpenStudy (anonymous):

Thanks